# thanks: Equal contributionthanks: Equal contribution

### v1

@@ II.4 Induction Gem
To preserve the divergence-free nature of the magnetic field, instead of directly evolving $\tilde{B}^{i}$ using the flux terms above for the induction equation, we evolve the vector potential $A_{k}$ using a staggered constrained-transport scheme [Balsara:1999, Etienne:2012, DelZanna_2003]. This guarantees that $\tilde{B}^{i}=\sqrt{\gamma}B^{i}$ remains divergence-free because, before each calculation of the evolution equations we compute
@@
To illustrate how the non-gauge terms in Eq. (3) are computed, let $\mathcal{E}_{i}\equiv-\epsilon_{ijk}v^{j}\tilde{B}^{k}$. Focusing on, e.g., the $z$-component and dropping the gauge terms, we find
@@
Having specified how the non-gauge flux term is computed, we now turn to the interpolation needed for the gauge-related terms and the scalar potential. As with the hydrodynamic RHS, $\tilde{\Phi}^{\mathrm{RHS}}$ and the gauge contribution to $A_{i}^{\mathrm{RHS}}$ require staggered variable quantities. However, we can use interpolation instead of reconstruction methods because gauge variables are not as sensitive to shocks. The Induction gem provides interpolator functions to facilitate the calculation of these terms.
The primary difficulty with these quantities is that they are all sampled at different cell locations, which complicates the interpolation scheme. The hydrodynamic quantities are cell-centered, and we denote these gridpoints as $(i,j,k)$, $(i+1,j,k)$, etc. The vector potential components are edge-centered—e.g., $A_{x}$ is evaluated at the point (i,j+1/2,k+1/2)(i,j+1/2,k+1/2)—and the scalar potential is vertex-centered. This staggering is summarized in Table 2.
@@
This dependency on the metric leads to several choices in how to compute the RHS. The original origIGM code uses the BSSN metric, but it can be more convenient in some codes to use the ADM metric. Additionally, different infrastructures assume the spacetime has different centerings. For example, the Carpet driver in the Einstein Toolkit assumes all variables have the same centering, so the spacetime is implicitly cell-centered (i.e., it is on the same grid as the hydrodynamic variables), while CarpetX supports different grid types and has vertex-centered spacetime variables but cell-centered hydrodynamic variables. Thus, we currently provide interpolators using the ADM metric with cell- and vertex-centered variables, and the BSSN metric with cell-centered variables.
@@ II.5 Reconstruction Gem
To use the HLL [Harten_1983] flux solver described in Sec. II.3, the primitive variables must be interpolated to cell faces without introducing oscillatory behavior. The Reconstruction gem provides several methods to perform this interpolation. The library offers minmod, monotonized-central, and superbee reconstruction, which apply total variation diminishing (TVD) linear interpolations [TVD] to the primitives. We refer to these methods collectively as the piecewise linear methods (PLM), noting that they are at best second-order accurate for smooth underlying solutions and revert to first-order at shocks.
GRHayL also provides the piecewise parabolic method (PPM) [PPM1, PPM2], which is third-order accurate in smooth regions and first-order accurate at extrema. We note that, other than PPM, all the reconstruction functions have the same argument lists, making switching between methods very simple. The PPM scheme requires additional inputs for its special handling of shocks and thus does not use the same interface. GRHayL initializes all PPM parameters to standard values by default, but the user can tune these parameters to achieve different behavior.
@@ II.6 Neutrinos Gem
GRHayL adopts what is perhaps the most popular approach for modeling neutrino physics in GRMHD simulations: a leakage scheme [vanRiper:1981mko, Ruffert:1995fs, Rosswog:2002rt, Rosswog:2003rv, Sekiguchi:2010ep, Sekiguchi:2011zd]. In this type of scheme, experimental data are used to parametrize analytic formulas for the neutrino emission and cooling rates in terms of the optical depths and opacities, resulting in a computationally inexpensive algorithm [OConnor:2009iuz, Ott:2012kr, Neilsen:2014hha, Radice:2016dwd, Siegel:2017jug, Endrizzi:2019trv, Murguia-Berthier:2021tnt, Werneck_IGM]. However, this scheme neglects the absorption of neutrinos and fails to account for heating and lepton number changes within hot ejecta from compact object mergers [Just:2021cls]. To account for these effects, more sophisticated techniques, like radiation transport with an M1 closure [1981MNRAS.194..439T, Shibata:2011kx] (see also [Richers:2020ntq, Radice:2021jtw]) or Monte Carlo methods [Miller:2019gig, Miller:2019dpt, Foucart:2020qjb, Foucart:2021mcb, Foucart:2021ikp], are necessary.
The implementation of the leakage scheme in GRHayL is provided by NRPyLeakage [Werneck_IGM]. [^3] We note that the neutrino source term for $\tau$ shown in Eq. (21) of [Werneck_IGM] is incorrect, as it is missing a factor of the lapse function. Our earlier implementation used the form $\sqrt{\gamma}\alpha u^{0}\mathcal{Q}=\sqrt{\gamma}W\mathcal{Q}$, but the correct contribution is $\alpha\sqrt{\gamma}W\mathcal{Q}$. This typo was also present in our implementation of the source term, but has since been fixed. Following [Ruffert:1995fs, Galeazzi:2013mia, Siegel:2017jug, Murguia-Berthier:2021tnt, Burrows:2004vq], we consider neutrino production from $\beta$-processes (electron/positron absorption by protons/neutrons), electron-positron pair annihilation, transverse plasmon decay, and nucleon-nucleon bremsstrahlung. The inverse reactions contribute to the total neutrino transport opacities, with inverse $\beta$-processes providing the dominant contribution to the optical depths of electron neutrinos and neutral-current scattering off neutrons the dominant contribution to the optical depths of heavy-lepton neutrinos.
For computing the optical depths, we adopt the general-purpose method of Nielsen et al. [Neilsen:2014hha] (see also [Siegel:2017jug, Murguia-Berthier:2021tnt, Werneck_IGM]), in which neutrinos leave the system by following the path of least resistance. Refer to Werneck et al. [Werneck_IGM] for more details on NRPyLeakage.
@@ III.1 Code Simplification
Over time, repeated inheritance of code has led to the build-up of technical debt within origIGM. While refactoring the code to write GRHayL, we found several places where the code could be simplified or condensed, with most of these improvements affecting the conservative-to-primitive routine and its surrounding functions.
@@ III.2 EoS Feature Extensions
We have also made several improvements and extensions to the latest version of IllinoisGRMHD in the Einstein Toolkit. The remaining subsections describe improvements to the evolution codes that implement GRHayL, and not the core library itself. This section details the extensions to IllinoisGRMHD’s EoS support. These include full support for hybrid polytropic EoSs, infrastructure to incorporate tabulated EoS and neutrino leakage into the official release version of origIGM, and the addition of the simple EoS.
origIGM originally only supported simple polytropic hybrid EoSs. While it had much of the code needed to extend to piecewise polytropic EoSs, elements of the code contained implicit assumptions regarding the EoS. Recent improvements have added tabulated EoS [Werneck_IGM] in a forked version of origIGM. With GRHayL, we fully incorporate the piecewise polytropic hybrid EoS feature into the release version of origIGM in the Einstein Toolkit, and we are working on tabulated EoS support. We also add the extra features necessary to use the simple EoS. While this is of less interest for physical simulations, the commonly used Balsara tests [Balsara_2001] assume this EoS, so the original origIGM code could not run those tests with the correct EoS for comparison with the exact solution.
@@ III.4 Con2Prim Improvements
The most significant changes and improvements to pre-existing code are in the Con2Prim implementation. origIGM uses the Noble2D routine, with the Font1D routine as the only backup. While Font1D virtually guarantees a successful inversion, it does so by setting $P=P_{\rm cold}$, which introduces spurious cooling in the system. This means that any Con2Prim failures near the surface of the stars will cause discontinuities in the pressure due to the thermal component of the pressure suddenly disappearing wherever Font1D is applied, limiting the code’s accuracy.
To address these issues, IllinoisGRMHD supports up to four Con2Prim routines of the user’s choosing. This design allows users to employ any of GRHayL’s Con2Prim methods and control their ordering at runtime. For hybrid EoS with entropy advection evolution, a suggested combination is Noble2D, Palenzuela1D, Noble1D_entropy, and finally Palenzuela1D_entropy.
@@
This multi-stage approach is designed to improve the likelihood that Con2Prim returns valid primitive variables. It also avoids the use of Font1D and atmospheric reset for as long as possible so as to not discard any thermal contributions. Note that Font1D is currently not optional in IllinoisGRMHD, requiring the user to change the code in order to disable it.
@@ III.6 Memory Usage
The first major change is the removal of all BSSN grid variables from the origIGM thorn. The original code used both ADM and BSSN grid variables, often recomputing them in several places. We eliminated dependencies on BSSN grid variables, greatly simplifying the GRHayL code and removing non-essential grid variables from IllinoisGRMHD. We also improved the efficiency of the hydrodynamic RHS computations in GRHayL by generating optimized code via NRPy, which removed the need for additional grid variables to store $T^{\mu\nu}$.
@@
All these changes result in a memory overhead reduction of over 40% from origIGM to IllinoisGRMHD. While entropy evolution and experimental tabulated EoS support do require a few additional grid variables, they still introduce far fewer variables than the number of variables that we have removed. The new GRHayLHD GRHD thorn has no magnetic variables, so it reduces the memory usage by $\sim$80% compared to origIGM. These changes represent a significant reduction of memory usage and access.
@@ IV.1 Einstein Toolkit
Before discussing our numerical results, we first describe in more detail the infrastructures in which we have implemented GRMHD codes using GRHayL. The Einstein Toolkit [ET_2023_11] is a collection of codes for astrophysics simulations. Most of these codes are built on the Cactus framework [Cactus], which provides a flexible infrastructure that allows many different code modules (called “thorns”) to be compiled together. The origIGM code is one such thorn. While Cactus provides many features, the actual numerical grids, parallelism, and related low-level details are implemented by driver thorns. The driver currently used for most production-level simulations is the Carpet driver, which provides AMR, implements MPI parallelism, and handles memory allocation during the simulation.
Similarly, CarpetX is a newly released driver that builds on the AMReX framework [amrex]. One major change with this driver is the automation of ghost-zone synchronization and boundary-condition application. This feature was added to Cactus in the PreSync update [Cupp_2018, Cupp_2019] but had not been implemented in many Carpet-based codes, including origIGM. CarpetX is designed with this new functionality in mind, and CarpetX-compatible codes are expected to provide the option for this more advanced inter-processor communication paradigm. Combined with the new looping and grid-function syntax, origIGM would need to be changed significantly to work with CarpetX. As such, an infrastructure-agnostic solution is highly desirable to reduce the work overhead of maintaining code for both drivers.
@@
Similarly, we provide a CarpetX-based thorn named GRHayLHDX, which again depends on GRHayLib. The IllinoisGRMHDX thorn is also in development, with expected inclusion in a future release of the Einstein Toolkit. With the inclusion of CarpetX in the Einstein Toolkit, providing reliable GRMHD codes for the new driver is critical for future science using the Einstein Toolkit, with AsterX [Kalinani:2024rbk] a noteworthy example. One important caveat is that CarpetX also supports GPUs. While work to support GPUs has begun, GRHayL currently only supports CPU code. As such, the CarpetX thorns run on the host at the moment. Once the library functions are available on the device, the thorns’ code can be easily switched to GPU-based computation.
@@ IV.2 BlackHoles@Home
BlackHoles@Home is an infrastructure built on the code-generation package NRPy [Ruchlin:2018, Etienne_nrpytutorial]. NRPy allows for efficient generation of optimized C code from symbolic Python expressions. BlackHoles@Home leverages this Python package to produce minimal but efficient numerical relativity code using high-order finite-differencing and time-stepping algorithms with extrapolation and high-order radiation outer boundary conditions. The infrastructure takes advantage of NRPy’s C code generation to produce state-of-the-art code for binary black hole simulations. The core goal of the infrastructure is to reduce the memory usage and inherent inefficiencies of box-in-box AMR on Cartesian grids by using multi-coordinate, multi-patch grids.
Previously, this infrastructure only supported vacuum spacetime evolutions. We have developed two GRHayL-based GR(M)HD codes for BlackHoles@Home: the GRHD code GRoovy and its GRMHD extension NRPyCartMHD. GRoovy [Jacques:2024pxh] implements GRHayL within BlackHoles@Home, yielding a GRHD code for fully dynamical spacetimes that supports single-patch curvilinear coordinates via a reference-metric formulation [Montero:2012yr, Baumgarte:2012xy, Mewes:2020vic]. NRPyCartMHD is a specialization of GRoovy that extends it to GRMHD while restricting the coordinate system to Cartesian coordinates. We are currently developing a supercomputer-ready version of GRoovy that leverages Charm++ task-based parallelism through superB [Jadoo:2025lgm].
@@ V Code Tests
We approach testing and validation through two complementary avenues. First, the core GRHayL library has many function-level tests that are regularly performed. The details of our continuous integration testing methods are given in Appendix B. Second, we compare physics simulations produced in several different infrastructures with results from origIGM. Through this testing, we validate the behavior of the replacements for origIGM by showing that the differences between the GRHayL-based codes and origIGM are within expectations. Additionally, we demonstrate successful parallel development of GRMHD codes using GRHayL in BlackHoles@Home, Carpet, and CarpetX.
@@
Balsara et al. [Balsara_2001] provide a very challenging battery of tests using one-dimensional GRMHD flows characterized by sharp features—along with extreme Lorentz factors or magnetization—that push GRMHD codes to their limits. These tests also have analytic solutions, allowing for direct validation of a numerical code’s results. While origIGM is capable of passing these tests, one must remove its pressure floors and ceilings and adjust some EoS functions in order to do so, preventing them from being run without manual changes to the code. Due to our changes and improvements, IllinoisGRMHD can run these tests without modification by using the simple EoS option. Here we briefly describe the Balsara tests and present results comparing IllinoisGRMHD and GRHayLHD with analytical results and GRHydro, another trusted GRMHD code in the Einstein Toolkit.
@@
As shown in the table, all of these tests have very small ratios of gas pressure to magnetic pressure. In addition, Balsara 3 has a pressure shock of four orders of magnitude without any change to the density, and the Balsara 4 test has an enormous Lorentz factor of over 22. Simply put, these tests all push an MHD code to the limits of the physics they are designed to model and serve as excellent first tests that an MHD code is working properly.
@@
One drawback to the Balsara tests is that they all have non-zero magnetic fields, which means that the analytical results cannot be used for purely hydrodynamic codes. However, we can still use a modified version of the Balsara tests to compare different codes. Since the shocks are still significant even in just the hydrodynamic variables, this still serves as an excellent test of the purely hydrodynamic GRHayL codes. We therefore compare and validate the hydrodynamic codes using a “Balsara 0” test, obtained from Balsara 1 by setting the magnetic fields to zero.
@@
The primary difference between the two codes is in the magnetic sector. GRHydro only supports evolving the magnetic field directly or evolving a cell-centered vector potential, while IllinoisGRMHD evolves a staggered (i.e. edge-centered) vector potential. We choose to compare against GRHydro with magnetic field evolution, which is the default and more heavily tested evolution scheme in the code. Further, it has been shown that evolution of the cell-centered vector potential leads to increased noise around shocks (see e.g. [Spritz]).
In addition to these numerical codes, we use the analytical solution produced by the code presented in [Giacomazzo_2006]. Fig. 3 shows the results for the Balsara tests with IllinoisGRMHD and GRHydro, overlaid with the exact solution. We see that the two codes agree very well not only with the exact solution, but also with each other. In the Balsara 3 test, IllinoisGRMHD captures the shock better due to using PPM instead of minmod.
V.2 TOV Neutron Star
To further validate GRHayL and its implementations within the Einstein Toolkit and the BlackHoles@Home infrastructures, we performed simulations with single neutron stars using Tolman–Oppenheimer–Volkoff (TOV) initial data. The TOV solution is a spherically symmetric solution to Einstein’s equations with a non-zero stress-energy tensor. In the following subsections we show numerical results from evolving TOV initial data in a fully dynamical, three-dimensional spacetime with no symmetry assumptions. All initial data used for these tests were generated using the TOV solver within NRPy.

### v2

@@ II.4 Induction Gem
To preserve the divergence-free nature of the magnetic field, instead of directly evolving $\tilde{B}^{i}$ using the flux terms above for the induction equation, we evolve the vector potential $A_{k}$ using a staggered constrained-transport scheme [21, 45, 36]. This guarantees that $\tilde{B}^{i}=\sqrt{\gamma}B^{i}$ remains divergence-free because, before each calculation of the evolution equations we compute
@@
To illustrate how the nongauge terms in Eq. (3) are computed, let $\mathcal{E}_{i}\equiv-\epsilon_{ijk}v^{j}\tilde{B}^{k}$. Focusing on, e.g., the $z$-component and dropping the gauge terms, we find
@@
Having specified how the nongauge flux term is computed, we now turn to the interpolation needed for the gauge-related terms and the scalar potential. As with the hydrodynamic RHS, $\tilde{\Phi}^{\mathrm{RHS}}$ and the gauge contribution to $A_{i}^{\mathrm{RHS}}$ require staggered variable quantities. However, we can use interpolation instead of reconstruction methods because gauge variables are not as sensitive to shocks. The Induction gem provides interpolator functions to facilitate the calculation of these terms.
The primary difficulty with these quantities is that they are all sampled at different cell locations, which complicates the interpolation scheme. The hydrodynamic quantities are cell centered, and we denote these gridpoints as $(i,j,k)$, $(i+1,j,k)$, etc. The vector potential components are edge centered—e.g., $A_{x}$ is evaluated at the point (i,j+1/2,k+1/2)(i,j+1/2,k+1/2)—and the scalar potential is vertex centered. This staggering is summarized in Table 2.
@@
This dependency on the metric leads to several choices in how to compute the RHS. The original origIGM code uses the Baumgarte-Shaapiro-Shibata-Nakamura (BSSN) metric, but it can be more convenient in some codes to use the ADM metric. Additionally, different infrastructures assume the spacetime has different centerings. For example, the Carpet driver in the Einstein Toolkit assumes all variables have the same centering, so the spacetime is implicitly cell centered (i.e., it is on the same grid as the hydrodynamic variables), while CarpetX supports different grid types and has vertex-centered spacetime variables but cell-centered hydrodynamic variables. Thus, we currently provide interpolators using the ADM metric with cell- and vertex-centered variables, and the BSSN metric with cell-centered variables.
@@ II.5 Reconstruction Gem
To use the HLL [61] flux solver described in Sec. II.3, the primitive variables must be interpolated to cell faces without introducing oscillatory behavior. The Reconstruction gem provides several methods to perform this interpolation. The library offers minmod, monotonized-central, and superbee reconstruction, which apply total variation diminishing linear interpolations [70] to the primitives. We refer to these methods collectively as the piecewise linear methods, noting that they are at best second-order accurate for smooth underlying solutions and revert to first order at shocks.
GRHayL also provides the piecewise parabolic method (PPM) [32, 72], which is third-order accurate in smooth regions and first order accurate at extrema. We note that, other than PPM, all the reconstruction functions have the same argument lists, making switching between methods very simple. The PPM scheme requires additional inputs for its special handling of shocks and thus does not use the same interface. GRHayL initializes all PPM parameters to standard values by default, but the user can tune these parameters to achieve different behavior.
@@ II.6 Neutrinos Gem
GRHayL adopts what is perhaps the most popular approach for modeling neutrino physics in GRMHD simulations: a leakage scheme [114, 101, 99, 98, 105, 104]. In this type of scheme, experimental data are used to parametrize analytic formulas for the neutrino emission and cooling rates in terms of the optical depths and opacities, resulting in a computationally inexpensive algorithm [88, 90, 84, 95, 108, 40, 83, 115]. However, this scheme neglects the absorption of neutrinos and fails to account for heating and lepton number changes within hot ejecta from compact-object mergers [66]. To account for these effects, more sophisticated techniques, like radiation transport with an M1 closure [111, 107] (see also [97, 94]) or Monte Carlo methods [79, 80, 52, 51, 53], are necessary.
The implementation of the leakage scheme in GRHayL is provided by NRPyLeakage [115]. [^3] We note that the neutrino source term for $\tau$ shown in Eq. (21) of [115] is incorrect, as it is missing a factor of the lapse function. Our earlier implementation used the form $\sqrt{\gamma}\alpha u^{0}\mathcal{Q}=\sqrt{\gamma}W\mathcal{Q}$, but the correct contribution is $\alpha\sqrt{\gamma}W\mathcal{Q}$. This typo was also present in our implementation of the source term, but has since been fixed. Following [101, 54, 108, 83, 26], we consider neutrino production from $\beta$-processes (electron/positron absorption by protons/neutrons), electron-positron pair annihilation, transverse plasmon decay, and nucleon-nucleon bremsstrahlung. The inverse reactions contribute to the total neutrino transport opacities, with inverse $\beta$-processes providing the dominant contribution to the optical depths of electron neutrinos and neutral-current scattering off neutrons the dominant contribution to the optical depths of heavy-lepton neutrinos.
For computing the optical depths, we adopt the general-purpose method of Nielsen et al. [84] (see also [108, 83, 115]), in which neutrinos leave the system by following the path of least resistance. Refer to Werneck et al. [115] for more details on NRPyLeakage.
@@ III.1 Code Simplification
Over time, repeated inheritance of code has led to the buildup of technical debt within origIGM. While refactoring the code to write GRHayL, we found several places where the code could be simplified or condensed, with most of these improvements affecting the conservative-to-primitive routine and its surrounding functions.
@@ III.2 EoS Feature Extensions
We have also made several improvements and extensions to the latest version of IllinoisGRMHD in the Einstein Toolkit. The remaining subsections describe improvements to the evolution codes that implement GRHayL, and not the core library itself. This section details the extensions to IllinoisGRMHD’s EoS support. These include full support for hybrid polytropic EoS, infrastructure to incorporate tabulated EoS and neutrino leakage into the official release version of origIGM, and the addition of the simple EoS.
origIGM originally only supported simple polytropic hybrid EoS. While it had much of the code needed to extend to piecewise polytropic EoS, elements of the code contained implicit assumptions regarding the EoS. Recent improvements have added tabulated EoS [115] in a forked version of origIGM. With GRHayL, we fully incorporate the piecewise polytropic hybrid EoS feature into the release version of origIGM in the Einstein Toolkit, and we are working on tabulated EoS support. We also add the extra features necessary to use the simple EoS. While this is of less interest for physical simulations, the commonly used Balsara tests [22] assume this EoS, so the original origIGM code could not run those tests with the correct EoS for comparison with the exact solution.
@@ III.4 Con2Prim Improvements
The most significant changes and improvements to preexisting code are in the Con2Prim implementation. origIGM uses the Noble2D routine, with the Font1D routine as the only backup. While Font1D virtually guarantees a successful inversion, it does so by setting $P=P_{\rm cold}$, which introduces spurious cooling in the system. This means that any Con2Prim failures near the surface of the stars will cause discontinuities in the pressure due to the thermal component of the pressure suddenly disappearing wherever Font1D is applied, limiting the code’s accuracy.
To address these issues, IllinoisGRMHD supports up to four Con2Prim routines of the user’s choosing. This design allows users to employ any of GRHayL’s Con2Prim methods and control their ordering at run-time. For hybrid EoS with entropy advection evolution, a suggested combination is Noble2D, Palenzuela1D, Noble1D_entropy, and finally Palenzuela1D_entropy.
@@
This multistage approach is designed to improve the likelihood that Con2Prim returns valid primitive variables. It also avoids the use of Font1D and atmospheric reset for as long as possible so as to not discard any thermal contributions. Note that Font1D is currently not optional in IllinoisGRMHD, requiring the user to change the code in order to disable it.
@@ III.6 Memory Usage
The first major change is the removal of all BSSN grid variables from the origIGM thorn. The original code used both ADM and BSSN grid variables, often recomputing them in several places. We eliminated dependencies on BSSN grid variables, greatly simplifying the GRHayL code and removing nonessential grid variables from IllinoisGRMHD. We also improved the efficiency of the hydrodynamic RHS computations in GRHayL by generating optimized code via NRPy, which removed the need for additional grid variables to store $T^{\mu\nu}$.
@@
[…]
@@ IV.1 Einstein Toolkit
Before discussing our numerical results, we first describe in more detail the infrastructures in which we have implemented GRMHD codes using GRHayL. The Einstein Toolkit [102] is a collection of codes for astrophysics simulations. Most of these codes are built on the Cactus framework [58], which provides a flexible infrastructure that allows many different code modules (called “thorns”) to be compiled together. The origIGM code is one such thorn. While Cactus provides many features, the actual numerical grids, parallelism, and related low-level details are implemented by driver thorns. The driver currently used for most production-level simulations is the Carpet driver, which provides AMR, implements MPI parallelism, and handles memory allocation during the simulation.
Similarly, CarpetX is a newly released driver that builds on the AMReX framework [116]. One major change with this driver is the automation of ghost-zone synchronization and boundary condition application. This feature was added to Cactus in the PreSync update [34, 33] but had not been implemented in many Carpet-based codes, including origIGM. CarpetX is designed with this new functionality in mind, and CarpetX-compatible codes are expected to provide the option for this more advanced interprocessor communication paradigm. Combined with the new looping and grid-function syntax, origIGM would need to be changed significantly to work with CarpetX. As such, an infrastructure-agnostic solution is highly desirable to reduce the work overhead of maintaining code for both drivers.
@@
Similarly, we provide a CarpetX-based thorn named GRHayLHDX, which again depends on GRHayLib. The IllinoisGRMHDX thorn is also in development, with expected inclusion in a future release of the Einstein Toolkit. With the inclusion of CarpetX in the Einstein Toolkit, providing reliable GRMHD codes for the new driver is critical for future science using the Einstein Toolkit, with AsterX [67] a noteworthy example. One important caveat is that CarpetX also supports GPUs. While work to support GPUs has begun, GRHayL currently only supports CPU code. As such, the CarpetX thorns run on the host at the moment. Once the library functions are available on the device, the thorns’ code can be easily switched to GPU-based computation.
@@ IV.2 BlackHoles@Home
BlackHoles@Home is an infrastructure built on the code-generation package NRPy [100, 46]. NRPy allows for efficient generation of optimized C code from symbolic Python expressions. BlackHoles@Home leverages this Python package to produce minimal but efficient numerical relativity code using high-order finite-differencing and time-stepping algorithms with extrapolation and high-order radiation outer boundary conditions. The infrastructure takes advantage of NRPy’s C code generation to produce state-of-the-art code for binary black hole simulations. The core goal of the infrastructure is to reduce the memory usage and inherent inefficiencies of box-in-box AMR on Cartesian grids by using multicoordinate, multipatch grids.
Previously, this infrastructure only supported vacuum spacetime evolutions. We have developed two GRHayL-based GR(M)HD codes for BlackHoles@Home: the GRHD code GRoovy and its GRMHD extension NRPyCartMHD. GRoovy [63] implements GRHayL within BlackHoles@Home, yielding a GRHD code for fully dynamical spacetimes that supports single-patch curvilinear coordinates via a reference-metric formulation [81, 24, 74]. NRPyCartMHD is a specialization of GRoovy that extends it to GRMHD while restricting the coordinate system to Cartesian coordinates. We are currently developing a supercomputer-ready version of GRoovy that leverages Charm++ task-based parallelism through superB [64].
@@ V Code Tests
We approach testing and validation through two complementary avenues. First, the core GRHayL library has many function-level tests that are regularly performed. The details of our continuous-integration testing methods are given in Appendix B. Second, we compare physics simulations produced in several different infrastructures with results from origIGM. Through this testing, we validate the behavior of the replacements for origIGM by showing that the differences between the GRHayL-based codes and origIGM are within expectations. Additionally, we demonstrate successful parallel development of GRMHD codes using GRHayL in BlackHoles@Home, Carpet, and CarpetX.
@@
Balsara et al. [22] provide a very challenging battery of tests using one-dimensional GRMHD flows characterized by sharp features—along with extreme Lorentz factors or magnetization—that push GRMHD codes to their limits. These tests also have analytic solutions, allowing for direct validation of a numerical code’s results. While origIGM is capable of passing these tests, one must remove its pressure floors and ceilings and adjust some EoS functions in order to do so, preventing them from being run without manual changes to the code. Because of our changes and improvements, IllinoisGRMHD can run these tests without modification by using the simple EoS option. Here we briefly describe the Balsara tests and present results comparing IllinoisGRMHD and GRHayLHD with analytical results and GRHydro, another trusted GRMHD code in the Einstein Toolkit.
@@
As shown in the table, all of these tests have very small ratios of gas pressure to magnetic pressure. In addition, Balsara 3 has a pressure shock of 4 orders of magnitude without any change to the density, and the Balsara 4 test has an enormous Lorentz factor of over 22. Simply put, these tests all push an MHD code to the limits of the physics they are designed to model and serve as excellent first tests that an MHD code is working properly.
@@
One drawback to the Balsara tests is that they all have nonzero magnetic fields, which means that the analytical results cannot be used for purely hydrodynamic codes. However, we can still use a modified version of the Balsara tests to compare different codes. Since the shocks are still significant even in just the hydrodynamic variables, this still serves as an excellent test of the purely hydrodynamic GRHayL codes. We therefore compare and validate the hydrodynamic codes using a “Balsara 0” test, obtained from Balsara 1 by setting the magnetic fields to zero.
@@
The primary difference between the two codes is in the magnetic sector. GRHydro only supports evolving the magnetic field directly or evolving a cell-centered vector potential, while IllinoisGRMHD evolves a staggered (i.e., edge-centered) vector potential. We choose to compare against GRHydro with magnetic field evolution, which is the default and more heavily tested evolution scheme in the code. Further, it has been shown that evolution of the cell-centered vector potential leads to increased noise around shocks (see, e.g., [30]).
In addition to these numerical codes, we use the analytical solution produced by the code presented in [56]. Figure 3 shows the results for the Balsara tests with IllinoisGRMHD and GRHydro, overlaid with the exact solution. We see that the two codes agree very well not only with the exact solution, but also with each other. In the Balsara 3 test, IllinoisGRMHD captures the shock better due to using PPM instead of minmod.
V.2 Tolman-Oppenheimer-Volkoff (TOV) Neutron Star
To further validate GRHayL and its implementations within the Einstein Toolkit and the BlackHoles@Home infrastructures, we performed simulations with single neutron stars using TOV initial data. The TOV solution is a spherically symmetric solution to Einstein’s equations with a nonzero stress-energy tensor. In the following subsections we show numerical results from evolving TOV initial data in a fully dynamical, three-dimensional spacetime with no symmetry assumptions. All initial data used for these tests were generated using the TOV solver within NRPy.